OpenStudy (anonymous):

what is the slope of a line that passes thrught the origin and the point (-6,2)

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

-1/3

OpenStudy (anonymous):

but how do u figure that out

OpenStudy (anonymous):

Well the equation of a line is y-y1=m(x-x1) where y1 and x1 are the y and x coordinates and m is the slope. The equation of the line that passes through origin (0,0) is y=mx and equation through y-2=m(x+6), which simplifies to y=mx+6m+2. These are 2 simultaneous equations. y=mx and y=mx+6m+2 Subtract them from each other, gives: [y=mx y=mx+6m+2] 0=0+6m+2 so we are left with 6m+2=0 rearrange to find m. 6m=-2 m=-2/6 =-1/3.

OpenStudy (anonymous):

the equations are both y=-1/3 x when the m=-1/3 is substituted.Check for yourself. Hope this helps

OpenStudy (anonymous):

thx soo much!!!